Exercise-15: Linear system of equations-Eigenvalues 215 In Exercises through 9, solve the system X = AX. 2. A= 4 4 - 8
Answer #1

4 4 8 eigenvalues (A-1)=0 4-s ng -8- 2) 4 (4-1) (-8-1) +36-0 -32+ 8x - 41+2+36=0 A taxta=0 (1+2)=0 (A= -2 conosponding eigen(3) now -g GP, - 9 h = 3 GP1 = 3+9 P2 P = 3 +9P2 6 V2 = 3 +9P2 let B= P2 N2 = 1) now the solution can be at tv, tev given as.